OpenStudy (anonymous):

\[sin^{-1}(1)\] \[sin(1)=\frac\pi 2\] that would give us \[(1,\frac \pi 2)\] the inverse would be \[(\frac \pi 2,1)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or is it the other way around....kinda confusing

OpenStudy (anonymous):

well not really what you wrote you just should write \[\sin^{-1}(1)=\frac{\pi}{2}\]

OpenStudy (anonymous):

you want \(\sin^{-1}(1)\) that is \(\frac{\pi}{2}\) because \(\sin(\frac{\pi}{2})=1\)

OpenStudy (anonymous):

yep makes sense. Thanks again!